# Switch-JustDance: Benchmarking Whole-Body Motion Tracking Policies Using a Commercial Console Game

### v2

@@ 5.3 Candidate Controllers
    GMT [chen2025gmt]: learns a unified motion-tracking policy by leveraging adaptive sampling and a mixture-of-experts architecture for diverse whole-body motions. For Smo, we do not expose future frames and use a constant hold-last-frame assumption over the lookahead. For Dyn, future reference frames are available and included in the observation stack.
@@
    TWIST [ze2025twist]: leverages human motion capture to train a general robot motion controller for teleoperation tasks.
@@
    Any2Track [zhang2025any2track]: learns a tracking policy using a two-stage RL framework that augments a general tracker with online disturbance adaptation to handle varied real-world conditions.
@@ 6 Benchmark Results
The results indicate that current humanoid controllers are incomparable to human-level performance. A human expert can reach the theoretical maximum score of the game, and participants in our user study achieve an average score of 8564. In comparison, even the best-performing controller reaches only about one-third of the human baseline score (Any2Track 2900 vs. human 8546). Interestingly, when comparing human motion against all controllers in the Real-Dyn setting, we find that these controllers actually achieve lower MPJPE than human players, with the best MPJPE reaching 88.8 mm while human players get 112.7 mm. This suggests that the robots can track the reference pose more closely at the frame level than human participants. However, this improved accuracy comes at the cost of stability: robots with lower MPJPE also exhibit lower success rates. Furthermore, human players produce much smoother movements, showing lower jerk and acceleration compared to all controllers.
@@ Appendix A Implementation Details for Real-Time Streaming of Smoothed gameplay (Smo)
Appendix B Song Selection for Benchmark Evaluation
@@
Applying these criteria, we select five songs for our benchmark, as summarized in Table B1.
@@
Table B1: Summary of selected Just Dance tracks with game editions, difficulty, and durations. (lvl #) is the in-game difficulty. we group lvl 1–2 as Easy and 3–4 as Hard.
Appendix C IRB and Human Subjects Approval

### v3

@@ 5.3 Candidate Controllers
    GMT [5]: learns a unified motion-tracking policy by leveraging adaptive sampling and a mixture-of-experts architecture for diverse whole-body motions. For Smo, we do not expose future frames and use a constant hold-last-frame assumption over the lookahead. For Dyn, future reference frames are available and included in the observation stack.
@@
    TWIST [39]: leverages human motion capture to train a general robot motion controller for teleoperation tasks.
@@
    Any2Track [41]: learns a tracking policy using a two-stage RL framework that augments a general tracker with online disturbance adaptation to handle varied real-world conditions.
@@ 6 Benchmark Results
The results indicate that current humanoid controllers are incomparable to human-level performance. A human expert can reach the theoretical maximum score of the game, and participants in our user study achieve an average score of 8538. In comparison, even the best-performing controller reaches only about one-third of the human baseline score (Any2Track 2900 vs. human 8538). Interestingly, when comparing human motion against all controllers in the Real-Dyn setting, we find that these controllers actually achieve lower MPJPE than human players, with the best MPJPE reaching 88.8 mm while human players get 111.9 mm. This suggests that the robots can track the reference pose more closely at the frame level than human participants. However, this improved accuracy comes at the cost of stability: robots with lower MPJPE also exhibit lower success rates. Furthermore, human players produce much smoother movements, showing lower jerk and acceleration compared to all controllers.
@@ Appendix A Implementation Details for Real-Time Streaming of Smoothed gameplay (Smo)
Appendix B Broader Embodiment and Generalization
By design, our framework is agnostic to different robot morphologies. In terms of scoring, the Just Dance system is designed for a global population and evaluates motion across diverse body sizes and proportions, making it largely morphology-agnostic. For retargeting, our pipeline builds on the open-sourced GMR framework, which has been successfully applied to multiple humanoid morphologies, including Booster and H1. To further validate generality, we apply our pipeline to a HOVER [15]-controlled H1 robot in simulation, where the robot successfully follows the reference dance trajectory. A representative snapshot is shown in Fig. B1.
Figure B1: Snapshot of our pipeline being applied to a HOVER [15]-controlled H1 robot in simulation, dancing to Old Town Road; from 0:35 to 0:39.
Appendix C Song Selection for Benchmark Evaluation
@@
Applying these criteria, we select five songs for our benchmark, as summarized in Table C1.
@@
Table C1: Summary of selected Just Dance tracks with game editions, difficulty, and durations. (lvl #) is the in-game difficulty. we group lvl 1–2 as Easy and 3–4 as Hard.
Appendix D IRB and Human Subjects Approval